# Structural Anisotropy Stabilises Asymmetric Beating in Instability Driven Filaments

### v1

@@ Abstract
Asymmetries and anisotropies are widespread in biological systems, including in the structure and dynamics of cilia and eukaryotic flagella. These microscopic, hair-like appendages exhibit asymmetric beating patterns that break time-reversal symmetry needed to facilitate fluid transport at the cellular level. The intrinsic anisotropies in ciliary structure can promote preferential beating directions, further influencing their dynamics. In this study, we employ numerical simulation and bifurcation analysis of a mathematical model of a filament driven by a follower force at its tip to explore how intrinsic curvature and direction-dependent bending moduli impact filament dynamics. Our results show that while intrinsic curvature is indeed able to induce asymmetric beating patterns when filament motion is restricted to a plane, this beating is unstable to out of plane perturbations. Furthermore, we find that a 3D whirling state seen for isotropic filament dynamics can be suppressed when sufficient asymmetry or anisotropy are introduced. Finally, for bending moduli ratios as low as 2, we demonstrate that combining structural anisotropy with intrinsic curvature can stabilise asymmetric beating patterns, highlighting the crucial role of anisotropy in ciliary dynamics.
@@ 1 Introduction
Asymmetry is ubiquitous in nature across multiple scales, from the chirality of molecules [1] to the left-right asymmetry in the placement of internal organs in the human body [2]. Many animals utilise asymmetry for mating or survival, such as flatfish like flounders and halibut, which have both eyes on one side of the head to allow them to lie flat against the ocean floor and avoid predators [3]. Anisotropy, here referring to the direction dependence of mechanical response of a material, is also common in nature, a simple example being the human hand, which is easier to bend in one direction than in any other direction. Fibrous materials such as wood are anisotropic due to the natural grain formed by fibre alignment, meaning the strength and flexibility of these materials are direction-dependent [4].
@@
Cilia and eukaryotic flagella are hairlike appendages which exhibit both asymmetric and anisotropic properties. These filaments are responsible for driving fluid flow at the microscale [5]. They can be found on the surfaces of swimming cells and microorganisms such as Volvox [6], collectively beating to allow the algae to navigate through flows [7, 8], or lining epithelial cells in our own bodies to facilitate fluid transport, such as mucus in the airways [9], or cerebrospinal fluid in the brain [10, 11, 12]. Symmetry breaking is essential for fluid transport at this scale [13]. Accordingly, cilia often undergo asymmetric beating oscillations, to ensure net fluid motion and cellular transport. Respiratory cilia for example, beat with distinct fast, low-curvature effective strokes, and high-curvature, slower, recovery strokes [14, 15]. Nodal cilia in the embryo, on the other hand, undergo tilted whirling motions [16, 17, 18], that direct fluid flow to the left and hence breaking left-right symmetry in the embryo.
@@
Such asymmetries in ciliary dynamics are thought to be the result of the internal structure of cilia known as the axoneme. While the configuration of the axoneme can vary [19], motile cilia most commonly have a 9+2 axonemal structure [5, 20, 21], consisting of a central pair of microtubules, surrounded by 9 pairs of microtubules, called microtubule doublets. The microtubule doublets are held relative to each other by structural proteins called nexin. The bridge between one pair of doublets is known to be much stiffer than the others [5, 22], indicating that this pair have limited motion relative to each other. The orientation of the central pair and the stiffer bridge suggest a plane in which it is easier for the axoneme to bend. It has been observed that the beat plane for cilia aligns with this plane [5], suggesting that the direction of the cilium beat could be a result of anisotropy in axonemal structure. In previous numerical studies modelling cilia, this has been included explicitly by considering a stiffer bridge between this pair of doublets [23], or implicitly by increasing the perpendicular bending modulus [24]. It has also often been used as motivation for restricting dynamics to be entirely planar [25].
@@
While these modelling choices can reproduce planar beating, the resulting dynamics are often symmetric and do not reproduce differences in the effective and recovery strokes observed experimentally. Additional asymmetry must be considered to generate more accurate simulations which match experiments. Some works achieve this by incorporating constraints on material properties of the filament, such as curvature-dependent bending moduli [23] in which the bending modulus can take different values depending on the local curvature of the filament’s centerline, or through biased kinetics of dynein motors [26, 27, 28]. A widely accepted approach, however, includes introducing an intrinsic curvature to the filament [26, 27, 29, 30, 31], in which it is assumed that the rest state of the cilium has finite curvature. These numerical studies are successful in demonstrating realistic cilia beats, with a close resemblance to the beats observed in nature, but are often restricted to two dimensional dynamics. Further studies are necessary to determine whether these dynamics persist when the filament is free to deform in any direction.
@@
In this paper we explore the effect of asymmetries and anisotropies in a 3D active filament model. To drive filament oscillations, we utilise the widely-studied follower force model [25, 32, 33, 34, 35, 36, 37, 38], whereby a single compressive force is applied to the tip of the filament, against the tangent. While the follower force model does not represent the sliding-mechanism known to generate ciliary beating, this model has been shown to exhibit beating and whirling dynamics for sufficient forcing values, which are reminiscent of cilia beating patterns observed in nature. As such, the follower force model has been used explore instabilities underpinning cilia dynamics [25, 39], understand the influence of fluid rheology on cilia’s fluid-pumping ability [31, 40], study filament beat coordination and collective dynamics [41], and model biofilaments in cytoplasmic streaming in Drosophila oocytes [42] and actin-myosin dynamics [43]. In the next section, we describe the active filament model we utilise to generate simulations, and the computational tools we employ to find steady and periodic states, and to analyse their stability. Next, we include asymmetry through a non-trivial intrinsic curvature to an otherwise-isotropic filament, and establish the dynamics in both 2D, matching those observed in the literature, and 3D. We find that the asymmetric planar beating found in 2D is inherently unstable in the full 3D model, and that the whirling solution can be eliminated entirely if the imposed asymmetry is large enough. Following this, we discuss the effect of direction-dependent bending moduli in the absence of intrinsic curvature, showing how whirling can again be eradicated if the anisotropy in the bending moduli is sufficiently large. Finally, we investigate how this anisotropy can be exploited on filaments with intrinsic curvature to recover and stabilise asymmetric beating dynamics in 3D. In particular, we show that bending moduli ratios as low as 2 are sufficient to stabilise asymmetric beating, a value agreeing with structural anisotropies that have been measured for cilia.
@@ 2 Numerical methods
Figure 1: Schematics of the model. (a) We consider a filament of length $L$ and radius $a$, clamped at its base $(s=0)$ to an infinite no-slip planar wall, and driven by a follower force imposed at the free end $(s=L).$ (b) The bending modulus anisotropy and intrinsic curvature are defined through the bending moment, which acts on the filament cross-section. It is expressed in terms of the filament’s material frame, $(\hat{\bm{t}},\hat{\bm{\mu}},\hat{\bm{\nu}})$. (c) The intrinsic curvature, $\kappa_{0}=L\tilde{\kappa_{0}}$ is defined to be the rest-curvature of the filament under zero forcing.
The methods employed here largely follow those described in [44, 37], and we provide a brief description below for completeness.
@@ 2.1 Filament Model
The filament has length $L$ and cross-sectional radius $a$, and is clamped at its based to a no-slip planar surface. The filament is forced at its free end by a follower force, a compressive tangential force, as shown in Figure 1a. The filament centerline, $\bm{Y}(s,t),$ is parameterised by arc-length, $s\in[0,L]$, and evolves over time, $t\in[0,\infty)$. Filament deformation is described by the orthonormal local basis $\{\bm{\hat{t}}(s,t),\bm{\hat{\mu}}(s,t),\bm{\hat{\nu}}(s,t)\}$. For the filament dynamics, we use the framework outlined in [44]. Due to the small length scales and velocity scales associated to ciliary motion, the Reynolds number is small and so we neglect the inertia of the fluid and the filament. Consequently, force and moment balances along the filament length yield
@@
and the internal moments are given by the constitutive relation [44],
@@
where $\tilde{\kappa}_{0}=\kappa_{0}/L$ defines the intrinsic curvature, such that the rest configuration of the filament is a planar arc with constant curvature $\kappa_{0},$ as sketched in Figure 1c. Therefore, larger values of $\kappa_{0}$ correspond to a higher curvature rest configurations for the filament. Also defined in Eq. (4) are $K_{B}$, the filament bending modulus in the direction of intrinsic curvature, $\beta K_{B}$, the bending modulus in the perpendicular plane, and $K_{T}$, the twisting modulus. Choosing $\beta>1$ leads to preferred bending in the $\hat{\bm{\nu}}-$direction, and $\beta<1$ leads to preferred bending in the $\hat{\bm{\mu}}-$direction. Correspondingly, setting $\beta\rightarrow\infty$ should restrict the filament dynamics to the $(\hat{\bm{\nu}},\hat{\bm{t}})$-plane entirely. Choosing $\beta=1,\kappa_{0}=0$ corresponds to the isotropic, symmetric case considered in [37].
The base of the filament $(s=0)$ is clamped to a no-slip planar surface, fixing the position and local orientation of the base to be $\bm{Y}(0,t)=\bm{0}$ and $(\hat{\bm{t}}(0,t),\hat{\bm{\mu}}(0,t),\hat{\bm{\nu}}(0,t))=(\hat{\bm{z}},\hat{\bm{y}},\hat{\bm{x}})$. The distal end $(s=L)$ is moment-free, $\bm{M}(L,t)=\bm{0}$, and is driven by the follower force,
@@
where $f$ is the non-dimensional follower force strength. We note that an alternative definition of the non-dimensional follower force can be reached by balancing with the elastic force in the $\hat{\bm{\nu}}$-direction. This would lead to a rescaled non-dimensional follower force, $\tilde{f}=\beta f$.
We discretise the filament into $N$ segments of length $\Delta L$, such that segment $i$ has centre $\bm{Y}_{i}$ and an orientation defined by the local frame, $\{\bm{\hat{t}}_{i},\bm{\hat{\mu}}_{i},\bm{\hat{\nu}}_{i}\}$ for $i=1,...,N$. We discretise Eqs (1), (2) and (3), using central differencing and rearrange to obtain
@@
for segment $i=1,...,N$, where $\bm{T}_{i}^{E}=\bm{M}_{i+1/2}-\bm{M}_{i-1/2}$ is the elastic torque, $\bm{F}_{i}^{C}=\bm{\Lambda}_{i+1/2}-\bm{\Lambda}_{i-1/2}$ and $\bm{T}_{i}^{C}=(\Delta L/2)\hat{\bm{t}}_{i}\times(\bm{\Lambda}_{i+1/2}+\bm{\Lambda}_{i-1/2})$ are the constraint forces and torques respectively, and $\bm{F}_{i}^{H}=\Delta L\bm{f}_{i}^{H}$ and $\bm{T}_{i}^{H}=\Delta L\bm{\tau}_{i}^{H}$ are the hydrodynamic force and torque respectively. The internal forces, $\bm{\Lambda}_{i+1/2},$ act as Lagrange multipliers enforcing the constraint, (8). The internal moments, $\bm{M}_{i+1/2}$, are defined by a discretised version of the constitutive law given in Eq (4).
@@
where $\mathcal{M}$ is the $6N\times 6N$ mobility matrix, $\bm{V}$ and $\bm{\Omega}$ are $6N\times 1$ vectors containing the translational and angular velocities of all segments respectively, and $\bm{F}^{H}$, $\bm{T}^{H}$ are $6N\times 1$ vectors containing the hydrodynamic forces and torques of all segments, respectively, which are defined through the force and moment balances, Eqs (6) and (7). In this work, we use the Rotne-Prager-Yamakawa (RPY) tensor [45], adapted to include the no-slip condition at $z=0$ [46], to define the mobility matrix, $\mathcal{M}$. Here, each segment is modelled as a particle with hydrodynamic radius equal to the cross-sectional radius of the filament, $a$.
After solving for the translational and rotational velocities of each segment, we integrate the system forwards in time. To understand the evolution of segment orientations, we utilise unit quaternions, $\bm{q}=(q_{0},q_{1},q_{2},q_{3})$ with $||\bm{q}||=1$ [44, 47]. Segment $i$ is defined to have a corresponding unit quaternion, $\bm{q}_{i}$, which maps the standard basis, $(\hat{\bm{x}},\hat{\bm{y}},\hat{\bm{z}})$, to the segment’s local basis at time $t,$ $(\hat{\bm{t}}_{i}(t)\hat{\bm{\mu}}_{i}(t)\hat{\bm{\nu}}_{i}(t))$, through a rotation matrix, $\bm{R}(\bm{q})$ (see [44] for further details). The evolution of each quaternion is described by
@@
where $\bullet$ is the quaternion product, defined for two quaternions $\bm{p}=(p_{0},\tilde{\bm{p}})$ and $\bm{q}=(q_{0},\tilde{\bm{q}})$ as $\bm{p}\bullet\bm{q}=(p_{0}q_{0}-\tilde{\bm{p}}\cdot\tilde{\bm{q}},\;p_{0}\tilde{\bm{q}}+q_{0}\tilde{\bm{p}}+\tilde{\bm{p}}\times\tilde{\bm{q}})$.
@@
we can discretise our system temporally using a second-order, geometric backward differential formula scheme, as described in [44]. These discretised equations, as well as the constraint given by (8), define a nonlinear system which we can solve iteratively using Broyden’s method [48].
In this work, the filament is discretised into $N=20$ segments of length $\Delta L=2.2a$, such that the filament has length $L=N\Delta L=44a.$ The twist modulus is taken to be equal to the bending modulus in the $\hat{\bm{\mu}}-$direction, $K_{B}=K_{T}$. In the isotropic case, where $\beta=1$ and $\kappa_{0}=0$, the magnitude of the twist has little effect on the dynamics [37]. We investigate the effect of varying three non-dimensional parameters; the intrinsic curvature, $\kappa_{0}$, as in (4); the bending anisotropy, quantified by $\beta,$ as in (4); and the nondimensional follower force, $f,$ as defined in (5). We focus on ranges of $\kappa_{0}$ and $\beta$ that follow values explored in other studies [24, 26, 27, 28, 29, 30, 31, 49], and a range of $f$ which allows us to access and explore the whirling and beating states. All initial value problems (IVPs) use the initial condition of a vertically upright filament, with random perturbations to the force on each segment. All timescales are non-dimensionalised by the filament’s relaxation time, $\tau$, the characteristic timescale associated with a bent filament returning to its undeformed configuration, which we obtain numerically as described in [37].

### v2

@@ Abstract
Asymmetries and anisotropies are widespread in biological systems, including in the structure and dynamics of cilia and eukaryotic flagella. These microscopic, hair-like appendages exhibit asymmetric beating patterns that break time-reversal symmetry needed to facilitate fluid transport at the cellular level. The intrinsic anisotropies in ciliary structure can promote preferential beating directions, further influencing their dynamics. In this study, we employ numerical simulation and bifurcation analysis of a mathematical model of a filament driven by a follower force at its tip to explore how intrinsic curvature and direction-dependent bending stiffnesses impact filament dynamics. Our results show that while intrinsic curvature is indeed able to induce asymmetric beating patterns when filament motion is restricted to a plane, this beating is unstable to out of plane perturbations. Furthermore, we find that a 3D whirling state seen for isotropic filament dynamics can be suppressed when sufficient asymmetry or anisotropy are introduced. Finally, for bending stiffness ratios as low as 2, we demonstrate that combining structural anisotropy with intrinsic curvature can stabilise asymmetric beating patterns, highlighting the crucial role of anisotropy in ciliary dynamics.
@@ 1 Introduction
Asymmetry is ubiquitous in nature across multiple scales, from the chirality of molecules (Garbacz2024IntroductionChirality, ) to the left-right asymmetry in the placement of internal organs in the human body (Hirokawa2009Left-rightFlow., ). Many animals utilise asymmetry for mating or survival, such as flatfish like flounders and halibut, which have both eyes on one side of the head to allow them to lie flat against the ocean floor and avoid predators (Palmer2009AnimalAsymmetry, ). Anisotropy, here referring to the direction dependence of mechanical response of a material, is also common in nature, a simple example being the human hand, which is easier to bend in one direction than in any other direction. Fibrous materials such as wood are anisotropic due to the natural grain formed by fibre alignment, meaning the strength and flexibility of these materials are direction-dependent (Kretschmann2010ChapterWood, ).
@@
Cilia and eukaryotic flagella are hairlike appendages which exhibit both asymmetric and anisotropic properties. These filaments are responsible for driving fluid flow at the microscale (Gibbons1981CiliaEukaryotes, ). They can be found on the surfaces of swimming cells and microorganisms such as Volvox (Pedley2016SquirmersSwimming, ), collectively beating to allow the algae to navigate through flows (Suarez2006SpermTract, ; Bennett2015AChlamydomonas, ), or lining epithelial cells in our own bodies to facilitate fluid transport, such as mucus in the airways (Smith2008ModellingClearance, ), or cerebrospinal fluid in the brain (Worthington1966CILIARYSURFACES, ; Guldbrandsen2014In-depthCSF-PR., ; Zhang2015AProteome, ). Symmetry breaking is essential for fluid transport at this scale (Purcell1977LifeNumber, ). Accordingly, cilia often undergo asymmetric beating oscillations, to ensure net fluid motion and cellular transport. Respiratory cilia for example, beat with distinct fast, low-curvature effective strokes, and high-curvature, slower, recovery strokes (Chilvers2000AnalysisMethods, ; Jorissen2023CiliaTransport, ). Nodal cilia in the embryo, on the other hand, undergo tilted whirling motions (Nonaka1998RandomizationProtein, ; Smith2008FluidCilia, ; Smith2011MathematicalCilia, ), that direct fluid flow to the left and hence breaking left-right symmetry in the embryo.
@@
Such asymmetries in ciliary dynamics are thought to be the result of the internal structure of cilia known as the axoneme. While the configuration of the axoneme can vary (Chaaban2017APolymers, ), motile cilia most commonly have a 9+2 axonemal structure (Gibbons1981CiliaEukaryotes, ; Marshall2006Cilia:Antenna, ; Gilpin2020TheFlagella, ), consisting of a central pair of microtubules, surrounded by 9 pairs of microtubules, called microtubule doublets. The microtubule doublets are held relative to each other by structural proteins called nexin. The bridge between one pair of doublets is known to be much stiffer than the others (Gibbons1981CiliaEukaryotes, ; AFZELIUS1959ElectronFixative., ), indicating that this pair have limited motion relative to each other. The orientation of the central pair and the stiffer bridge suggest a plane in which it is easier for the axoneme to bend. It has been observed that the beat plane for cilia aligns with this plane (Gibbons1981CiliaEukaryotes, ), suggesting that the direction of the cilium beat could be a result of anisotropy in axonemal structure. In previous numerical studies modelling cilia, this has been included explicitly by considering a stiffer bridge between this pair of doublets (Han2018SpontaneousCilia, ), or implicitly by increasing the perpendicular bending stiffness (Rallabandi2022Self-sustainedFlagellum, ). It has also often been used as motivation for restricting dynamics to be entirely planar (Bayly2016SteadyFlagella, ).
@@
While these modelling choices can reproduce planar beating, the resulting dynamics are often symmetric and do not reproduce differences in the effective and recovery strokes observed experimentally. Additional asymmetry must be considered to generate more accurate simulations which match experiments. Some works achieve this by incorporating constraints on material properties of the filament, such as curvature-dependent bending stiffness (Han2018SpontaneousCilia, ) in which the bending stiffness can take different values depending on the local curvature of the filament’s centerline, or through biased kinetics of dynein motors (Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Chakrabarti2022ACilia, ). A widely accepted approach, however, includes introducing an intrinsic curvature to the filament (Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Sartori2016DynamicFlagella, ; Cass2023TheFlagella, ; Wang2023Generalized-NewtonianFilament, ), in which it is assumed that the rest state of the cilium has finite curvature. These numerical studies are successful in demonstrating realistic cilia beats, with a close resemblance to the beats observed in nature, but are often restricted to two dimensional dynamics. Further studies are necessary to determine whether these dynamics persist when the filament is free to deform in any direction.
@@
In this paper we explore the effect of asymmetries and anisotropies in a 3D active filament model. To drive filament oscillations, we utilise the widely-studied follower force model (Bayly2016SteadyFlagella, ; DeCanio2017, ; Ling2018Instability-drivenMicrofilaments, ; Fily2020BucklingFilaments, ; Fatehiboroujeni2018NonlinearDrag, ; Fatehiboroujeni2021Three-dimensionalFlipping, ; Clarke2024BifurcationsFilaments, ; Schnitzer2025OnsetFilament, ), whereby a single compressive force is applied to the tip of the filament, against the tangent. While the follower force model does not represent the sliding-mechanism known to generate ciliary beating, this model has been shown to exhibit beating and whirling dynamics for sufficient forcing values, which are reminiscent of cilia beating patterns observed in nature. As such, the follower force model has been used to explore instabilities underpinning cilia dynamics (Bayly2016SteadyFlagella, ; Woodhams2022GenerationModels, ), understand the influence of fluid rheology on cilia’s fluid-pumping ability (Wang2023Generalized-NewtonianFilament, ; Link2024EffectFilament, ), study filament beat coordination and collective dynamics (Westwood2021CoordinatedSurfaces, ), and model biofilaments in cytoplasmic streaming in Drosophila oocytes (Stein2021SwirlingCytoskeleton, ) and actin-myosin dynamics (Sekimoto1995SymmetrySystem, ). In the next section, we describe the active filament model we utilise to generate simulations, and the computational tools we employ to find steady and periodic states, and to analyse their stability. Next, we include asymmetry through a non-trivial intrinsic curvature to an otherwise-isotropic filament, and establish the dynamics in both 2D, matching those observed in the literature, and 3D. We find that the asymmetric planar beating found in 2D is inherently unstable in the full 3D model, and that the whirling solution can be eliminated entirely if the imposed asymmetry is large enough. Following this, we discuss the effect of direction-dependent bending stiffness in the absence of intrinsic curvature, showing how whirling can again be eradicated if the anisotropy in the bending stiffnesses is sufficiently large. Finally, we investigate how this anisotropy can be exploited on filaments with intrinsic curvature to recover and stabilise asymmetric beating dynamics in 3D. In particular, we show that bending stiffness ratios as low as 2 are sufficient to stabilise asymmetric beating, a value agreeing with structural anisotropies that have been suggested for cilia (Ishijima1994FlexuralFlagella, ; Lindemann2016FunctionalFlagellum, ; Rallabandi2022Self-sustainedFlagellum, ).
@@ 2 Numerical methods
[…]
The methods employed here largely follow those described in (Scholler2021MethodsFilaments, ; Clarke2024BifurcationsFilaments, ), and we provide a brief description below for completeness.
@@ 2.1 Filament Model
The filament has length $L$ and cross-sectional radius $a$, and is clamped at its base to a no-slip planar surface. The filament is forced at its free end by a follower force, a compressive tangential force, as shown in Figure 1a. The filament centerline, $\bm{Y}(s,t),$ is parameterised by arc-length, $s\in[0,L]$, and evolves over time, $t\in[0,\infty)$. Filament deformation is described by the orthonormal local basis $\{\bm{\hat{t}}(s,t),\bm{\hat{\mu}}(s,t),\bm{\hat{\nu}}(s,t)\}$. For the filament dynamics, we use the framework outlined in (Scholler2021MethodsFilaments, ). Due to the small length scales and velocity scales associated to ciliary motion, the Reynolds number is small and so we neglect the inertia of the fluid and the filament. Consequently, force and moment balances along the filament length yield
@@
and the internal moments are given by the constitutive relation (Scholler2021MethodsFilaments, ),
@@
[…]
@@
where $f$ is the non-dimensional follower force strength. Through this non-dimensionalisation, the length is incorporated in the non-dimensional parameter, $f$. The impact of filament aspect ratio, $L/a$, in the symmetric and isotropic case ($\kappa_{0}=0$ and $\beta=1$) was investigated in (Clarke2024BifurcationsFilaments, ). We note that an alternative definition of the non-dimensional follower force can be reached by balancing with the elastic force in the $\hat{\bm{\nu}}-$direction. This would lead to a rescaled non-dimensional follower force, $\tilde{f}=\beta f$.
We discretise the filament into $N$ segments of length $\Delta L$, such that segment $i$ has centre $\bm{Y}_{i}$ and an orientation defined by the local frame, $\{\bm{\hat{t}}_{i},\bm{\hat{\mu}}_{i},\bm{\hat{\nu}}_{i}\}$ for i=1,,Ni=1,...,N. We discretise Eqs (1), (2) and (3), using central differencing and rearrange to obtain
@@
for i=1,,N1i=1,...,N-1. Here $\bm{T}_{i}^{E}=\bm{M}_{i+1/2}-\bm{M}_{i-1/2}$ is the elastic torque, $\bm{F}_{i}^{C}=\bm{\Lambda}_{i+1/2}-\bm{\Lambda}_{i-1/2}$ and $\bm{T}_{i}^{C}=(\Delta L/2)\hat{\bm{t}}_{i}\times(\bm{\Lambda}_{i+1/2}+\bm{\Lambda}_{i-1/2})$ are the constraint forces and torques respectively, and $\bm{F}_{i}^{H}=\Delta L\bm{f}_{i}^{H}$ and $\bm{T}_{i}^{H}=\Delta L\bm{\tau}_{i}^{H}$ are the hydrodynamic force and torque respectively. The internal forces, $\bm{\Lambda}_{i+1/2},$ act as Lagrange multipliers enforcing the constraint, (8). The internal moments, $\bm{M}_{i+1/2}$, are defined by a discretised version of the constitutive law given in Eq (4).
@@
where $\mathcal{M}$ is the $6N\times 6N$ mobility matrix, $\bm{V}$ and $\bm{\Omega}$ are $6N\times 1$ vectors containing the translational and angular velocities of all segments respectively, and $\bm{F}^{H}$, $\bm{T}^{H}$ are $6N\times 1$ vectors containing the hydrodynamic forces and torques of all segments, respectively, which are defined through the force and moment balances, Eqs (6) and (7). In this work, we use the Rotne-Prager-Yamakawa (RPY) tensor (Wajnryb2013GeneralizationTensors, ), adapted to include the no-slip condition at $z=0$ (Swan2007SimulationBoundary, ), to define the mobility matrix, $\mathcal{M}$. Here, each segment is modelled as a particle with hydrodynamic radius equal to the cross-sectional radius of the filament, $a$. Various hydrodynamic models have been used in this context. The follower force model appears to give rise to the same dynamics in the forcing ranges we consider, regardless of the details of the hydrodynamic model (Ling2018Instability-drivenMicrofilaments, ). For a comparison between RPY and nonlocal slender body theory, we direct the reader to (Maxian2021Integral-basedFlow, ).
[…]
@@
[…]
@@
we can discretise our system temporally using a second-order, geometric backward differential formula scheme, as described in (Scholler2021MethodsFilaments, ). These discretised equations, as well as the constraint given by (8), define a nonlinear system which we can solve iteratively using Broyden’s method (Broyden1965AEquations, ).
In this work, the filament is discretised into $N=20$ segments of length $\Delta L=2.2a$, such that the filament has length $L=N\Delta L=44a.$ This separation has been chosen to avoid overlap between neighbouring segments when the filament bends. The twist stiffness is taken to be equal to the bending stiffness in the $\hat{\bm{\mu}}-$direction, $K_{B}=K_{T}$. In the isotropic case, where $\beta=1$ and $\kappa_{0}=0$, the magnitude of the twist has little effect on the dynamics (Clarke2024BifurcationsFilaments, ). We investigate the effect of varying three non-dimensional parameters; the intrinsic curvature, $\kappa_{0}$, as in (4); the bending anisotropy, quantified by $\beta,$ as in (4); and the nondimensional follower force, $f,$ as defined in (5). We focus on ranges of $\kappa_{0}$ and $\beta$ that follow values explored in other studies (Rallabandi2022Self-sustainedFlagellum, ; Chakrabarti2019SpontaneousMicrofilaments, ; Chakrabarti2019HydrodynamicFilaments, ; Chakrabarti2022ACilia, ; Sartori2016DynamicFlagella, ; Cass2023TheFlagella, ; Wang2023Generalized-NewtonianFilament, ; Lindemann2016FunctionalFlagellum, ), and a range of $f$ which allows us to access and explore the whirling and beating states. All initial value problems (IVPs) use the initial condition of a vertically upright filament, with random perturbations to the force on each segment. All timescales are non-dimensionalised by the filament’s relaxation time, $\tau$, the characteristic timescale associated with a bent filament returning to its undeformed configuration, which we obtain numerically as described in (Clarke2024BifurcationsFilaments, ).